# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

### v1

@@ 3.2 Neural Field and Differentiable Rendering
Inspired by the success of UniPAD [53], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training scheme. Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
@@
where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [35]. Then we extract geometry features $\mathbf{f}_{\text{geo}}\in\mathbb{R}^{\hat{d}_{\text{geo}}}$ at $\mathbf{p}$ with $f^{\text{geo}}$, taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs,
$\mathbf{f}_{\text{geo}}=f^{\text{geo}}([\mathbf{p},\mathbf{f}]).$ (6)
With geometry features $\mathbf{f}_{\text{geo}}$, we predict the signed distance value $s\in\mathbb{R}$ [4, 28] and color value $s\in\mathbb{R}$ [44] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$. $f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron,
$s=f^{\text{SDF}}(\mathbf{f}_{\text{geo}})\ \ ,\ \ c=f^{\text{RGB}}(\mathbf{f}_{\text{geo}}).$ (7)
Differentiable Rendering. Similar to [32, 44], we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$. Next, we sample $N_{\text{ray}}$ points following [44] along each ray. Here each point along the ray can be interpreted by $\mathbf{p}=o+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$. Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$. Following [44], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
@@
Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $\mathcal{T}_{n}$ similar to [44]
$\mathcal{T}_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (9)
Based on $\mathcal{T}_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=\mathcal{T}_{n}\alpha_{n}$ [44] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
@@
$\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(\omega_{\text{L}}|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\omega_{\text{C}}\frac{1}{N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}|c_{i}-\tilde{c}_{i}|,$ (11)
where $\omega_{\text{L}}$, $\omega_{\text{sur}}$ and $\omega_{\text{C}}$ are weighting parameters to balance the contributions from the three losses. $s_{i}$ is the predicted signed distance value at the observed point along the sampled ray.
Figure 3: Inspiration of Curvature Sampling. From the two zoom-in areas, it can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature might provide more information for pre-training (objects like cars).
@@ 3.3 Curvature Sampling
In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$. Intuitively, uniform sampling can be used. But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling brings little improvement against separate pre-training, which is contradictory to our motivation. Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$. As shown in Figure 3, we are inspired by the observation that surface with higher curvature (surface of a vehicle) generally contains more information as compared to that with lower curvature (road plane). Therefore, we propose the Curvature Sampling for effective sampling. For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate the surface normal by deriving the signed distance function with respect to $\mathbf{p}$
@@
where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal. Then we normalized $\mathbf{n}$ to get the direction of the normal
$\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}.$ (13)
Here $||\cdot||_{2}$ is the L-2 norm. Next we estimate the mean curvature $\mathbf{c}\in\mathbb{R}^{3}$ at $\mathbf{p}$ by applying the vector differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$
$\mathbf{c}=\nabla_{\mathbf{p}}\tilde{\mathbf{n}}.$ (14)
For each point $\mathbf{p}_{n}$ in LiDAR point cloud, we compute the average of $\mathbf{c}_{n}$ as the sampling weights $\omega_{n}$ and sample $N_{\text{L}}$ LiDAR point cloud with a Multinomial Sampler implemented in PyTorch [35] for differentiable rendering. For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixel for $\mathcal{L}_{\text{rend}}$. As the estimation of curvature here is noisy especially during the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
3.4 Swapping Prototype Assignment Prediction
CLAP aims to utilize the interaction between camera and LiDAR encoders for joint unsupervised 3D representation learning. Meanwhile, we hope to learn a common feature space standing for segments of different objects in an unsupervised manner. Inspired by SwAV [3], we propose to use learnable prototypes to represent parts of the scenes. We first randomly initialize $N_{\text{K}}$ learnable prototypes $\mathbf{K}\in\mathbb{R}^{N_{\text{K}}\times d_{\text{K}}}$, each of which is a $d_{\text{K}}$ vectors. Then we pre-train the backbones to understand the interaction between different modalities with a swapping prototype assignment prediction loss [3]. Furthermore, we introduce an Expectation-Maximization scheme to optimize the learnable prototypes and a gram matrix minimization loss to avoid collapse [3] of the prototype learning.
Swapping Prototype Assignment Prediction. We project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
@@ 3.4 Swapping Prototype Assignment Prediction
where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$. We denote $N_{\text{3D}}=\hat{D}\times\hat{H}\times\hat{W}$ and then normalize and reshape the projected embeddings into $\dot{\mathbf{P}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$ and $\dot{\mathbf{I}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$. After that, similarity scores $\mathbf{S}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$ between 3D embeddings and prototypes are computed separately for LiDAR and camera branches
@@
Similar to SwAV [3], we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approach the double stochastic matrix in $N_{\text{sink}}$ iterations. We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$. The swapping prototype assignment loss is formulated with a temperature parameter $\tau$ [3],
@@
Expectation-Maximization. In order to guide the learnable prototypes to represent parts of the environment, we introduce the Expectation-Maximization algorithm [33] to further optimize the prototypes. In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$. Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is the same as minimizing the entropy of the similarity matrix. Hence, the EM loss is computed as,
@@
Table 1: Results for fine-tuning on 5% of training set in NuScenes [2]. “Init.” means the way to initialize models, including the baseline UniPAD [53]. To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [27]) until we observe convergence, which leads to the “Rand.” row. Then we pre-train the backbones with UniPAD [53] and the proposed CLAP, initialize BEVFusion with the pre-trained weights and then fine-tune the model with the same number iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font. We also indicate the performance improvement by green color. For detailed categories, we provide AP for them and “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone. All the results here are in %.
@@ 4 Experiments
Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks. In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [2] to demonstrate the effectiveness of CLAP. To begin with, we describe experiment setups in Section 4.1. Next, we show and analyze main results in Section 4.2. Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.4.
@@ 4.1 Settings
Datasets. We use the popular autonomous driving dataset NuScenes [2] to evaluate the performance of CLAP. NuScenes [2] uses one roof LiDAR and six surrounding cameras to collect data. The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz. The frequency of camera capturing is 12Hz. [2] conducts the synchronization and provides paired data of LiDAR point cloud and camera images. The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore. Each scene lasts for around 20 seconds and there are a total of 5.5 hours data. Following convention practice from [2, 42], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes. Pre-trainings are conducted on the whole training set and downstream task is 3D object detection under few-shot setting, where we randomly sample $5\%$ labeled data in the training set to train the randomly initialized model and fine-tune the pre-trained ones. The overall mAP (mean accurate precisions) and NDS (NuScenes Detection Score) are reported, along with AP (accurate precisions) for different categories.
@@
Downstream 3D Object Detectors. We conduct experiments on one of the SOTA fusion 3D object detector called BEVFusion [27]. BEVFusion is implemented in the popular code repository for autonomous driving perception called OpenPCDet [42]. We use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [2] as the evaluation metrics. We follow a similar setting in [48] to gradually increase training iterations of the model without pre-trained weights until convergence is observed. Here convergence means further increasing training iterations will not improve the performance. Then the number of training iterations is fixed for fine-tuning pre-trained models. This setting avoid the case that pre-training only accelerate convergence and make sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
Baseline Pre-training Method for Fusion Perception. We incorporate UniPAD [53] as the baseline 3D unsupervised pre-training method. We utilize the official implementation from [53] to pre-train the backbones of BEVFusion [27]. As pre-trained image backbones are broadly used for image feature extraction, the implementation of [53] also use a pre-trained image backbone from [8] to initiate their pre-training and the training set for this backbone does not involve any data from NuScenes [2]. We adopt this practice and also for training-from-scratch model, we initialize the camera encoder with the same pre-trained weights from [8].
@@
Implementation Details of CLAP. The feature channels for LiDAR embeddings $\hat{\mathbf{P}}$, image embeddings $\hat{\mathbf{I}}$, fusion embeddings $\hat{\mathbf{F}}$ and swapping prototype assignment prediction are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$. Sampling number for LiDAR point cloud and camera pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$, where $N_{\text{cam}}=6$ for NuScenes dataset [2]. The number of sample points along each ray is $N_{\text{ray}}=96$. Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$. For swapping prototype assignment prediction, we set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$. The temperature for swapping prediction loss is $\tau=1.0$. The loss weighting parameters are implemented as $\omega_{\text{proto}}=1.0$, $\omega_{\text{L}}=2.0$, $\omega_{\text{sur}}=0.1$, $\omega_{\text{C}}=0.1$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$. We use a learning rate of 0.00005 with a cosine learning schedule for pre-training and use mask augmentation for CLAP with a masking rate of 0.9.

### v2

@@ 3.2 Neural Field and Differentiable Rendering
Inspired by the success of UniPAD [58], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training. Different from [58], an additional surface signed distance field loss is applied to better optimize the scene geometry. Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
@@
where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [37]. Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ [5, 31] and color value $c\in\mathbb{R}^{3}$ [49] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
$s=f^{\text{SDF}}([\mathbf{p},\mathbf{f}])\ \ ,\ \ c=f^{\text{RGB}}([\mathbf{p},\mathbf{f}]),$ (6)
where $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
Differentiable Rendering. Similar to [33, 49], we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$. Next, we sample $N_{\text{ray}}$ points following [49] along each ray. Here each point along the ray can be interpreted by $\mathbf{p}=\mathbf{o}+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$. Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{RGB}}$ and $f^{\text{SDF}}$. Following [49], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
@@
Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to [49]
$t_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (8)
Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ [49] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
@@
$\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\frac{\omega_{\text{C}}}{3\cdot N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}\sum_{j=1}^{3}|c_{i}^{j}-\tilde{c}_{i}^{j}|,$ (10)
where $\omega_{\text{sur}}$ and $\omega_{\text{C}}$ are weighting parameters for the losses. $r_{i}$ is the observed range along the $i^{th}$ sampled ray and $s_{i}$ is the predicted signed distance value at the observed points. $c_{i}^{j}$ is the value of $j^{th}$ channel in the image pixel, where $j=\{1,2,3\}$ corresponds to RGB channels.
Figure 3: Inspiration of Curvature Sampling. It can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature provide more information (objects like cars).
@@ 3.3 Curvature Sampling
In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$. Intuitively, uniform sampling with depth can be used, same as “Memory-friendly Ray Sampling” in [58]. But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), this sampling method brings little improvement against separate pre-training, which is contradictory to our motivation. Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$. As shown in Figure 3, we are inspired by the observation that surface with higher curvature (surface of a vehicle) generally contains more information as compared to that with lower curvature (road plane). Therefore, we propose the Curvature Sampling for effective sampling. For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate the surface normal by deriving the signed distance function with respect to $\mathbf{p}$
@@
where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal. Then we normalized $\mathbf{n}$ to get the direction of the normal $\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}$. Here $||\cdot||_{2}$ is the L-2 norm. Next we apply the differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$ and get $\mathbf{c}\in\mathbb{R}^{3}$.
$\mathbf{c}=\frac{\delta\tilde{\mathbf{n}}}{\delta\mathbf{p}}.$ (12)
Then we estimate the geodesic curvature [48] of each point $\mathbf{p}_{n}$ in LiDAR point cloud by computing the norm of $\mathbf{c}_{n}$ and use it as the sampling weights $\omega_{n}$, that is $\omega_{n}=||\mathbf{c}_{n}||_{2}$. With $\omega_{n}$, we sample $N_{\text{L}}$ points with a Multinomial Sampler implemented in PyTorch [37] for differentiable rendering. For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixels for $\mathcal{L}_{\text{rend}}$. As the estimation of curvature here is noisy especially in the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
3.4 Prototype Learning
Curvature Sampling allows for joint pre-training, prompting us to explore how camera and LiDAR data can be used to understand “objectness” or “object parts” in an unsupervised way. We use learnable prototypes to represent segments of 3D scenes and establish a shared feature space that connects the two modalities. Firstly, we randomly initialize $N_{\text{K}}$ learnable prototypes $\mathbf{K}\in\mathbb{R}^{N_{\text{K}}\times d_{\text{K}}}$, each of which is a $d_{\text{K}}$ vectors. Then an Expectation-Maximization training approach is proposed to maximize the similarity between these prototypes and 3D embeddings from the two modalities. To delve deeper into the interplay between image semantics and LiDAR geometry, we use the swapping prototype prediction loss. Finally, to avoid the prototype collapsing into one same vector, we introduce a Gram Matrix Regularization Loss.
Expectation-Maximization. In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization [35] training scheme to optimize the prototypes. We first project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
@@ 3.4 Prototype Learning
where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$. We denote $N_{\text{3D}}=\hat{D}\times\hat{H}\times\hat{W}$ and then normalize and reshape the projected embeddings into $\dot{\mathbf{P}}\in\mathbb{R}^{N_{\text{3D}}\times d_{\text{K}}}$ and $\dot{\mathbf{I}}\in\mathbb{R}^{N_{\text{3D}}\times d_{\text{K}}}$. After that, similarity scores $\mathbf{S}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$ between 3D embeddings and prototypes are computed separately for LiDAR and camera branches
@@
In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$. Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is the same as minimizing the entropy of the similarity matrix. Hence, the EM loss is computed as,
Swapping Prototype Prediction. To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approximate to double stochastic matrix in $N_{\text{sink}}$ iterations. We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$. The swapping prototype prediction loss is formulated with a temperature parameter $\tau$, inspired by [4],
@@
Table 1: Results for fine-tuning on 5% of training set in NuScenes [3]. “Init.” means the way to initialize models. To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [30]) until we observe convergence, which leads to the “Rand.” row. Then we pre-train the backbones with ALSO [2], OCC-MAE [34], SLidR [38], PPKT [28], UniPAD [58] and CLAP. We initialize BEVFusion with the pre-trained weights and fine-tune the model with the same number iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font. We also indicate the performance improvement by green color. “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone. All the results here are in %.
@@ 4 Experiments
Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks. In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [3] and Waymo [45] to demonstrate the effectiveness of CLAP. To begin with, we describe experiment setups in Section 4.1. Next, we show and analyze main results in Section 4.2. Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.4.
@@ 4.1 Settings
Datasets. We use the popular autonomous driving dataset NuScenes [3] and Waymo [45] to evaluate the performance of CLAP. NuScenes [3] uses one roof LiDAR and six surrounding cameras to collect data. The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz. The frequency of camera capturing is 12Hz. [3] conducts the synchronization and provides paired data of LiDAR point cloud and camera images. The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore. Each scene lasts for around 20 seconds and there are a total of 5.5 hours data. Following convention practice from [3, 46], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes. Waymo [45] uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images. Following the same practice in [45, 46], the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes). All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes [3] $5\%$ and Waymo [45] $1\%$).
@@
Downstream 3D Object Detectors. For NuScenes [3], we select the SOTA 3D object detector called BEVFusion [30] and for Waymo [45], we use CenterPoint [59]. Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet [46]. For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [3] for NuScenes [3] and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo [45]. We follow a similar setting in [14, 53] to gradually increase training iterations of the from-scratch model until convergence is observed. Here convergence means further increasing training iterations will not improve the performance. Then the number of training iterations is fixed for fine-tuning pre-trained models. This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
Baseline Pre-training Method for Fusion Perception. We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO [2], 2) occupancy masked autoencoder called Occupancy-MAE [34], 3) multi-modality methods including SLidR [38], PPKT [28] and UniPAD [58]. We use the official implementations to pre-train the backbones with same setting as CLAP.
@@
Implementation Details of CLAP. The feature channels for embeddings $\hat{\mathbf{P}}$, $\hat{\mathbf{I}}$, $\hat{\mathbf{F}}$ and prototypes $\mathbf{K}$ are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$. Sampling number for point cloud and pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$. The number of sample points along each ray is $N_{\text{ray}}=96$. Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$. We set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$. The temperature for swapping prediction loss is $\tau=1.0$. The loss weighting parameters are implemented as $\omega_{\text{r}}=2.0$, $\omega_{\text{proto}}=1.0$, $\omega_{\text{sur}}=0.05$, $\omega_{\text{C}}=0.05$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$. We use torch.auto_grad() [37] to implement the derivatives in the curvature estimation. For more details about pre-training and fine-tuning, please refer to Appendix 1.